DX Actuator Addition Testing

15:47
  • Waiting for telescope
  • Preparing to put 560 and 561 into no ICC.
  • 92 modes!!!!
  • Loaded test_202257_400.sav with only 102 modes.
  • 376 has high forces, putting in no ICC
  • I made a mistake, I put them in no pos, now I've put them into no ICC
  • set with 114 modes.
16:54
  • Loop Closed.
  • Took 560, 561, 376 out of no ICC.
17:00
  • Recovering from all WFS processes disconnecting.
  • Loop Closed
  • Gain: 1.23 0.35 0.50
  • Dataset 170908
  • flat: test_170908_400.sav
  • Loaded with 222 modes
17:11
  • Removing 561 from pos
  • Last Flat Loaded with 222 modes
  • Gains 1.01 0.30 0.50
  • Dataset 172101
  • flat: test_172101_400.sav
  • Loaded with 220 modes
17:22
  • put 561 back in, removed 560
  • Default with 110 modes
  • last flat: 219 modes
  • Gains: 101 0.40 0.34
  • Dataset: 173136
  • Flat: test_171136_400.sav
  • Re-applied with 222 modes
17:33
  • Removing 561 and 560, going to last, best state from yesterday.
  • best flat application from yesterday: test_195523_400.sav
  • Still applied with 114 modes?
  • Gains: 0.51 0.20 0.20
  • Dataset: 174300
  • Flat: test_174300_400.sav
  • re-applied with 258 modes
17:49
  • Bump at about 8:00 position, caused by two actuators which are gone from pos and curr, will try to put back in
  • 262 and 263 put back into current only.
  • Re-applied last flat with 321 modes!!!!!!!
  • Gains: 0.51 0.20 0.20
  • Dataset: 175725
  • Flat: test_175725_400.sav
  • re-applied with 326 modes!!!
17:59
  • Will put 262 and 263 back into position.
  • Re applied default with 630 modes?!?!?!?!?!
  • Last flat with 630 modes?!?!?!
  • There is a bite out at 8:00 Trying to get multiple flats to restore it.
  • Dataset: 181012
  • Flat: test_181012_400.sav
  • Looking better
  • gains 0.95 0.31 1.11, about what it was last time, which I forgot to write.
  • Dataset: 181347
  • flat: test_181347_400.sav
18:18
  • Can't get chunk back, will put in 262, save a flat, and take it out again, in hopes of getting it back.
  • Flat applied with 629 modes.
  • Gain 0.87 0.20 1.11
  • Datset: 182516
  • Flat: test_182516_400.sav
18:27
  • Take out 262, put in 263
  • Not going to work, can't get the light back, taking it out.
  • Both cause a current of around -0.8, when the flat is set. Well out of bounds.
18:37
  • Removing 58, 169, 573 from "no curr" list.
  • Flat applied with 110 modes.
  • Gains 0.86 0.28 0.20
  • Dataset: 184809
  • Flat: test_184809_400.sav
  • Reset with 270
18:50
  • Putting those three back onto "no curr" list
  • removing 260 from "no pos" list
  • It doesn't work, the program won't load.
18:57
  • Putting 262 and 263 back into position.
  • Dataset: 191747
  • flat: test_191747_400.sav
  • Re applied with 630 modes
  • new gains: 0.88 0.30 1.11
  • Dataset: 192255
  • Flat: test_192255_400.sav
  • new gains: 0.96 0.20 1.19
  • Dataset: 192618
  • Flat: test_192618_400.sav
  • re-applied with 630 modes
  • new gains: 0.96 0.29 1.11
  • Dataset: 192953
  • new flat: test_192953_400.sav
  • Getting Closer
  • New Gains: 0.80 0.29 1.11
  • Dataset: 193338
  • New Flat: test_193338_400.sav
  • Manual Gains! 1.00 0.50 0.60
  • Dataset: 193842
  • New Flat: test_193842_400.sav
  • Really Getting There Now.
  • Manual Gains: 1.00 0.50 0.65
  • Dataset: 194254
  • flat: test_194254_400.sav
  • Gains 0.97 0.30 1.10
  • Dataset: 194726
  • flat: test_194726_400.sav
19:50
  • I'm going to try and remove them from ICC, to see if that helps.
  • Sets with 630 modes
  • The last flat really doesn't look too bad.
20:25
  • dist_flao1_ts4_atm_s0.8_L040.0_v_15.0_ovfreq1000.00_sd1983.fits
  • Gain: 0.87 0.59 1.60
  • Dataset: 20170503_202340
  • Gain: 0.97 0.48 0.90
  • 202627
  • Gain 1.02 0.44 0.70
  • 202928
  • dist_flao1_ts4_atm_s1.2_L040.0_v15.0_ovfreq1000.00_sd3892.fits
  • Gain 0.86 0.61 1.29
  • Dataset: 203352
  • Gain 0.86 0.61 0.54
  • Dataset: 203730
  • Gain 0.96 0.61 0.69
  • Dataset: 203931
  • dist_flao1_ts4_atm_s1.5_L040.0_v15.0_ovfreq1000.00_sd3892.fits
  • Gain: 0.97 0.68 1.60
  • Dataset: 204505
  • Gain 0.97 0.68 0.50
  • Dataset: 204641
  • Gain: 0.97 0.68 0.65
  • Dataset: 204904
20:57
  • Put 58 and 169 in both current and pos. Removed 262 and 263 in pos, they were causing a dark hole that couldn't be fixed and the HO2 gain always went to 1.6
  • Sets with 630 modes
  • Gains: 0.97 0.30 1.11
  • Dataset: 210954
  • Flat: test_210954_400.flat
  • Sets with 630 modes
  • Gains 0.94 0.20 1.11
  • Dataset: 211407
  • flat: test_211407_400.sav
21:16
  • Quick test with 262, 267 out in current as well, there is still a bit of a problem with that part of the mirror.
  • sets 630 modes
  • Much better looking pupils.
  • Gains: 0.87 0.39 1.10
  • Dataset: 212156
  • flat: test_212156_400.sav
  • sets with 630 modes
  • Dataset 212543
  • flat: test_212543_400.sav
  • Sets 630 modes
  • Dataset: 213231
  • Flat: test_213231_400.sav
21:43
  • putting in 72, 145, 151, 218, 223, 239, 307, 357, 367, from "no pos" list.
  • set with 638
  • Dataset: 215127
  • Manual gains: 1.00 0.30 0.30
  • Gains 0.51 0.20 0.56
  • Dataset: 215619
  • test_215619_400.sav
  • dist_flao1_ts4_atm_s1.0_L040.0_v15.0_ovfreq1000.00_sd3892.fits
  • Autogain: 0.96 0.59 0.50
  • Dataset 220221
  • Gain 0.96 0.59 0.70
  • Dataset: 220513
  • Gain 0.96 0.59 0.60
  • Dataset: 220621
22:16
  • Removed 115 from No Current
  • Added 235 and 236 to no pos, they have no magnet, according to comment in no current list.
  • sets with 639 modes
  • Dataset: 222345
  • flat: test_222345_400.sav
-- %USERSIG{GregoryTaylor - 2017-05-03}%

Comments

No Comment
Topic revision: r3 - 03 May 2017, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback